OpenStudy (anonymous):

Simplify. You may use the properties to help: -6(-2p + 3q − 6s)

OpenStudy (anonymous):

-6(-2p + 3q − 6s) = (-6)*(-2)p + (-6)*(3)q - (-6)(-6)s Whatever this simplifies to that is your answer.

OpenStudy (anonymous):

omfg how do you even simplify that

OpenStudy (anonymous):

Do you know how to multiply sign numbers?

OpenStudy (anonymous):

Multiplying sign numbers: (-)*(+) = - (+)*(-) = - (-)*(-) = + (+)*(+)= +

OpenStudy (anonymous):

nope math is like a foreign language to me lol

OpenStudy (anonymous):

Ok, So what is (-6)*(-2)p = ? , multiply the numbers

OpenStudy (anonymous):

If you don't know how to multiply sign numbers use a calculator. (-6)*(-2)p = 12p

OpenStudy (anonymous):

(-6)*(-2)p + (-6)*(3)q - (-6)(-6)s = 12p + (-18)q - 36s = 12p -18q-36s is your answer. Suggestion: If you are struggling with this you definitely need to go over multiplication, addition, subtraction and division of sign numbers with your instructor. Because, this is one of the foundation principles of mathematics, if you don't understand this you will have difficulty understanding other concepts in math.